Calculate the isoelectric point of glutamic acid
1
Expert's answer
2020-12-09T12:51:40-0500

pI=(pKa1+pKa3)÷2

pKa1=2.19

pKa2=9.67

pKa3=4.25


pI=(2.19+4.25)÷2

=6.44÷2

=3.22